OpenStudy (anonymous):

Hi I have a question about working with exponents that are larger than a regular scientific calculator can handle -- it asks for 21.9-19.70/(3.59x10^-422 m3) 1/2 [Please note, the 1/2 is superscript) I had some help earlier but it was for rooting a problem and I'm just unsure how to proceed with this one.

OpenStudy (anonymous):

WellI know that you would have to subtract 21.9-19.70

OpenStudy (ranga):

The denominator is: (3.59x10^-422 m3)^1/2 which can be simplified as follows: (3.59x10^-422 m3)^1/2 = sqrt(3.59 x 10^-422) = sqrt(3.59) * sqrt(10^-422) = sqrt(3.59) * 10^-211

OpenStudy (ranga):

Is m a variable or meter cubed?

OpenStudy (anonymous):

it's meter cubed

OpenStudy (anonymous):

I cannot do a square root of 10^-422 on my calculator though

OpenStudy (anonymous):

I mean I cannot enter more than two digits of notation on my calculator -- does that make sense?

OpenStudy (ranga):

You don't need a calculator to take square root of 10^-422. (10^-422)^1/2 = 10^-211

OpenStudy (anonymous):

oh. my mistake, i obviously have no idea how to solve it :P

OpenStudy (ranga):

In the problem, is the denominator just under 19.7 or under the whole (21.9 - 19.7) ?

OpenStudy (anonymous):

the whole 21.9-19.7

OpenStudy (anonymous):

exactly how i wrote it

OpenStudy (ranga):

Then you should put a parenthesis around the numerator so it would be grouped together. The way you have written it, only 19.7 is divided by the denominator.

OpenStudy (anonymous):

sorry that was the way it was typed in our homework

OpenStudy (anonymous):

here can i draw it? give me 2 seconds

OpenStudy (ranga):

In math textbooks, they write "vertically" with the numerator on top with a line below it and the denominator at the bottom and so it will be clear the whole thing is divided by the denominator. But when writing "horizontally" with no up and down, you should use parenthesis to make it clear the terms that are together.

OpenStudy (anonymous):